If Sally can paint a house in 4 hours, and John can paint the same house in 6 hours, how long will it take for both of them to paint the same house together?
Answer by stanbon(57369)

If Sally can paint a house in 4 hours, and John can paint the same house in 6 hours, how long will it take for both of them to paint the same house together?
--------------------
Sally time = 4 hrs/job ; rate = 1/4 job/hr
---
John time = 6 hrs/job ; rate = 1/6 job/hr
---
Together time = x hrs/job ; rate = 1/x job/hr
--------
Equation:
rate + rate = together rate
1/4 + 1/6 = 1/x
---
Multiply thru by 12x to get:
3x + 2x = 12
5x = 12
x = 12/5 = 2hr 24min (time to do the job together)
---------------
Cheers,
Stan H.
=============